# From Tokens to Words: On the Inner Lexicon of LLMs

### v2

@@ 7 Conclusion
The ability of LLMs to comprehend and generate language relies on intricate internal processes, and understanding these mechanisms is crucial for improving model performance as well as making them more efficient. In this work, we unraveled the word detokenization process, shedding light on how models internally transform fragmented sub-word tokens into coherent word representations formed at the last token. Our results indicate that this mechanism manifests in early to middle layers, where models attempt to reconstruct words by mapping them to an inner lexicon using their FFN layers. We provided evidence this lexicon is more exhaustive than the tokenizer’s vocabulary, and could facilitate the model’s ability to recognize words even amidst noise.
@@
Acknowledgements
This work was supported in part by the Israel Science Foundation (grant no. 2045/21) and by NSF-BSF grant 2020793. In addition, we thank Iddo Yosha for his valuable feedback.
@@ Acknowledgements
To further analyze results from Section 3, we conduct a failure analysis, focusing on false negatives (FN) and false positives (FP). The former, valid words misclassified as gibberish, are often rare or complex words, suggesting that the model’s internal vocabulary may lack representations for less common words. On the other hand, false positives—where gibberish is misclassified as a word—typically involved nonwords that closely resemble valid words, likely due to shared sub-word structures. See Table 2 for a few examples.
(a) accuracy postfix and last token
(b) positive predictions over prefix and last token
Figure 6: Comparisons between last token and the prefixes. (Left) we see that the accuracy over the prefix tokens reaching only 61% vs the 89% with the last token’s representation. (Right) a more detailed statistics over the positive predictions. The TP rate over the last token’s representations reaching 84% while reaching only 43% for the prefix ones
@@
Table 2: Examples of false negative (FN) and false positive (FP) for the word vs. nonword experiment (Section 3).
@@ Appendix B Statistical Analysis of Token Aggregation in Multi-token Words
We repeat the multi-token experiments Section 5.2 for 3- and 4-token words. Results Fig. 7 show a very similar trend to the 2-token words (Fig. 5).
We present the detailed statistical analysis conducted of our experiments in Section 5.2 to examine token aggregation in multi-token words compared to single-token words (control group). The objective is to investigate whether the model exhibits significantly different attention patterns between these groups, indicative of the detokenization process.
@@
To determine significance, we performed two one-sided t-tests per layer for the Llama2-7B case: (1) Testing whether attention to prefix tokens in multi-token words is higher than to previous tokens in single-token words. (2) Testing whether attention to previous tokens in single-token words is higher than to prefix tokens in multi-token words.
Table 3 shows the p-values and significance levels for each layer. Significance levels are denoted as ns (not significant), and *** ($p<0.001$).
@@
The results reveal that in layers 1 and 2, attention to prefix tokens in multi-token words is significantly higher than in single-token words, suggesting the early phase of Token Aggregation. From Layers 3 to 17, attention to single-token words is higher, indicating a shift in attention focus from prefix attention in relation regular close tokens. Notably, there are intermittent increases in attention to prefix tokens at Layers 18, 21, 25, 27, 29, and 30, possibly signaling a transfer into a prediction ensembling phase in which the attention in general is less important (Artzy & Schwartz, 2024) and therefore we don’t see a coherent pattern.
In conclusion, the attention mechanisms differ significantly between multi-token and single-token words, confirming that the detokenization process involves initial amplification of attention to sub-word tokens followed by a reduction as the model processes word representations.
Appendix C Word Retrieval for Single-Token and Multi-Token Words Across All Models Comparison
This section presents a detailed comparison of word retrieval performance (Section 4) across several models (Llama2-7B, Llama3-8B, Yi-6B, and Mistral-7B) for both single-token and multi-token words. The evaluation focuses on how effectively each model retrieves the original word across layers, especially in challenging cases like artificially separated single-token words, typos, and multi-token words.
@@ Appendix C Word Retrieval for Single-Token and Multi-Token Words Across All Models Comparison
Across all experiments, we observe a similar trend where the retrieval performance increases over the first several layers, peaks around the middle layers, and then decreases in the later layers. The main difference across models lies in the peak performance, especially in cases involving typos and multi-token words, where more advanced models such as Llama3-8B and Mistral-7B demonstrate superior performance in reconstructing the original word representations.
Appendix D Cumulative Word Retrieval for Single-Token and Multi-Token Words
In this section, we present the results of cumulative word retrieval (Section 4.1) across various models, focusing on both single-token and multi-token words. For each model, we analyze the ability of the LLM to retrieve the original word from sub-word tokens across its layers. The cumulative retrieval is calculated as the proportion of words that are successfully retrieved at each layer, with the percentage increasing as more words are recovered throughout the model’s layers.
Single-Token Words: The cumulative word retrieval for single-token words—those that are artificially split into sub-word tokens (via typographical errors or manual splits)—shows a rapid increase in retrieval success in the early layers. For Llama2-7B, for instance, cumulative retrieval reaches 93.2% for words split by manual intervention, and 66% for words affected by typos by the middle layers. This pattern is observed across models, with retrieval generally peaking around layers 15-20.
Multi-Token Words: For multi-token words, which are naturally split due to being out-of-vocabulary for the tokenizer, the cumulative retrieval process follows a similar trajectory. However, in models like Llama2-7B, the retrieval peaks earlier in the model, with a cumulative retrieval rate of 77.41%. Other models like Llama3 and Yi show higher cumulative retrieval rates, suggesting improved efficiency in handling multi-token words, potentially due to larger model capacities and internal dictionaries.
@@ Appendix D Cumulative Word Retrieval for Single-Token and Multi-Token Words
Appendix E Introducing Typos for Single-Token Words
In this section, we describe the process of introducing typos into single-token words to split them into multiple tokens (Section 4.1). The modification applies to words longer than four characters and involves randomly performing one of three operations: substituting two characters, deleting a character, or inserting a new character. By introducing these slight variations, the word becomes unfamiliar to the tokenizer, causing it to be divided into multiple smaller tokens during tokenization. Particularly, this process results in splitting words into 2–5 tokens. Table 4 shows examples of the different splits.
Description perturbed . new tokens
@@ Appendix E Introducing Typos for Single-Token Words
Table 4: Examples of the different typos we consider, exemplified by perturbing the single-token word “development” (Section 4.1).

### v3

@@ 7 Conclusion
The ability of LLMs to comprehend and generate language relies on intricate internal processes, and understanding these mechanisms is crucial for improving model performance and efficiency. In this work, we unraveled the word detokenization process, shedding light on how models internally transform fragmented sub-word tokens into coherent word representations formed at the last token. Our results indicate that this mechanism manifests in early to middle layers, where models attempt to reconstruct words by mapping them to an inner lexicon using their FFN layers. We provided evidence this lexicon is more exhaustive than the tokenizer’s vocabulary, and could help models to recognize words even amidst noise.
@@
Acknowledgments
This work was supported in part by the Israel Science Foundation (grant no. 2045/21) and by NSF-BSF grant 2020793. We are grateful to Iddo Yosha for his valuable insights and thank the anonymous reviewers for their constructive feedback.
@@ Acknowledgments
To further analyze our results from Sec. 3, we conduct a failure analysis, focusing on false negatives (FN) and false positives (FP). Our analysis indicates that the former—valid words misclassified as gibberish—are often rare and complex words, suggesting that the model’s internal vocabulary may lack representations for infrequent words. On the other hand, false positives—where gibberish is misclassified as a word—typically involves nonwords that closely resemble valid words, likely due to shared sub-word structures. See Tab. 2 for a few examples.
@@
Table 2: Examples of false negative (FN) and false positive (FP) for the word vs. nonword experiment (Sec. 3).
We further conduct additional experiments to compare the performance of our custom nonword dataset with the linguistically-motivated ARC Nonword Database (Rastle et al., 2002), which contains morphologically plausible nonwords designed to resemble real words to humans. Our results (Fig. 7) indicate a higher differentiability on the ARC dataset compared to our own dataset, which highlights that our nonword creation procedure effectively mitigates potential biases. This result strengthens our conclusions: the last token provides significant cues for distinguishing words from nonwords, particularly in the middle layers of the model after several layers of processing.
Figure 7: Accuracy comparison by layer for distinguishing words from nonwords. The ARC Nonword Database (blue line) exhibits the highest accuracy across layers, potentially due to its morphologically plausible nonwords. In contrast, our dataset (orange line, last token) achieves slightly lower accuracy, demonstrating the effectiveness of our bias-mitigated nonword generation process. The penultimate token from our dataset (green line) shows significantly lower accuracy, highlighting the importance of the last token in distinguishing words from nonwords.
In another experiment, we evaluate the penultimate token representation against nonword tokens for words of length three or more tokens. For example, in a word like “unhappiness,” we use the inner token “h” (e.g., un-h-appiness). Unlike the final tokens, penultimate tokens are poorly distinguishable from nonwords, achieving significantly lower accuracies. These results suggest that the model’s ability to separate words from nonwords is highly dependent on the position and completeness of token representations.
Finally, we expand the comparison of our dataset across three additional models—Llama3-8B, Mistral-7B, and Yi-6B. Our results across these models (Fig. 8) mirror the trends observed with Llama2-7B, with accuracy improving as layers deepened but peaking and gradually declining in later layers. This consistency further supports our findings that the model’s ability to classify words vs. nonwords relies on nuanced patterns in internal token representations, unaffected by co-occurrence biases or morphological artifacts.
Figure 8: Mean true status accuracy by layer for distinguishing words from nonwords across multiple models (Llama2-7B, Llama3-8B, Mistral-7B, Yi-6B) using our dataset. All models exhibit similar trends: accuracy improves across initial layers, peaks in the middle layers, and declines in deeper layers. This pattern demonstrates the robustness of our findings across different model architectures and sizes.
@@ Appendix B Statistical Analysis of Token Aggregation in Multi-token Words
We repeat the multi-token experiments Sec. 5.2 for 3- and 4-token words. Our results (Fig. 9) show a very similar trend to 2-token words (Fig. 5).
We further present the detailed statistical analysis conducted of our experiments in Sec. 5.2 to examine token aggregation in multi-token words compared to single-token words (control group). The objective is to investigate whether the model exhibits significantly different attention patterns between these groups, indicative of the detokenization process.
@@
To determine significance, we perform two one-sided t-tests per layer for the Llama2-7B case: (1) Testing whether attention to prefix tokens in multi-token words is higher than to previous tokens in single-token words. (2) Testing whether attention to previous tokens in single-token words is higher than to prefix tokens in multi-token words. Table 3 shows the p-values and significance levels for each layer. Significance levels are denoted as ns (not significant), and *** ($p<0.001$).
@@
The results reveal that in layers 1 and 2, attention to prefix tokens in multi-token words is significantly higher than in single-token words, suggesting the early phase of Token Aggregation. From Layers 3 to 17, attention to single-token words is higher, indicating a shift in attention focus from prefix attention in relation regular close tokens. Notably, there are intermittent increases in attention to prefix tokens at Layers 18, 21, 25, 27, 29, and 30, possibly signaling a transfer into a prediction ensembling phase in which the attention in general is less important (Ben Artzy & Schwartz, 2024) and therefore we don’t see a coherent pattern.
In conclusion, the attention mechanism differs significantly between multi-token and single-token words, indicating that the detokenization process involves initial amplification of attention to sub-word tokens followed by a reduction as the model obtains whole-word representations.
Appendix C Word Retrieval for Single-Token and Multi-Token Words Across Models
This section presents a detailed comparison of word retrieval performance (Sec. 4) across several models (Llama2-7B, Llama3-8B, Yi-6B, and Mistral-7B) for both single-token and multi-token words. The evaluation focuses on how effectively each model retrieves the original word across layers, especially in challenging cases like artificially separated single-token words, typos, and multi-token words.
Across all experiments (Fig. 10), we observe a similar trend where the retrieval rate increases over the first several layers, peaks around the middle layers, and then decreases in the later layers. The main difference across models lies in the peak performance, especially in cases involving typos and multi-token words, where more advanced models such as Llama3-8B and Mistral-7B demonstrate superior performance in reconstructing the original word representations.
@@ Appendix C Word Retrieval for Single-Token and Multi-Token Words Across Models
C.1 Cumulative Word Retrieval for Single-Token and Multi-Token Words
Fig. 11 shows the results of cumulative word retrieval (Sec. 4) across various models, focusing on both single-token and multi-token words. For each model, we analyze the ability of the LLM to retrieve the original word from sub-word tokens across its layers. The cumulative retrieval is calculated as the proportion of words that are successfully retrieved at each layer, with the percentage increasing as more words are recovered throughout the model’s layers.
Single-token words
The cumulative word retrieval for single-token words—those that are artificially split into sub-word tokens (via typographical errors or manual splits)—shows a rapid increase in retrieval success in the early layers. For Llama2-7B, for instance, cumulative retrieval reaches 93.2% for words split by manual intervention, and 66% for words affected by typos by the middle layers. This pattern is observed across models, with retrieval generally peaking around layers 15-20.
Multi-token words
For multi-token words, which are naturally split due to being out-of-vocabulary for the tokenizer, the cumulative retrieval process follows a similar trajectory. However, in models like Llama2-7B, the retrieval peaks earlier in the model, with a cumulative retrieval rate of 77.41%. Other models like Llama3 and Yi show higher cumulative retrieval rates, suggesting improved efficiency in handling multi-token words, potentially due to larger model capacities and internal dictionaries.
@@ C.1 Cumulative Word Retrieval for Single-Token and Multi-Token Words
Appendix D Introducing Typos for Single-Token Words
In this section, we describe the process of introducing typos into single-token words to split them into multiple tokens (Sec. 4.1). The modification applies to words longer than four characters and involves randomly performing one of three operations: substituting two characters, deleting a character, or inserting a new character. By introducing these slight variations, the word becomes unfamiliar to the tokenizer, causing it to be divided into multiple smaller tokens during tokenization. Particularly, this process results in splitting words into 2–5 tokens. Table 4 shows examples of the different splits.
Description perturbed new tokens
@@ Appendix D Introducing Typos for Single-Token Words
Table 4: Examples of the different typos we consider, exemplified by perturbing the single-token word “development” (Sec. 4.1).
Appendix E Ablation Experiment on Suffix-split Words
In Sec. 5, to test the role of FFN layers in detokenization, we conduct an intervention-based experiment measuring how ablations to FFN updates affect word retrieval. We run this experiment on all single-token words from WikiText-103 that end with one of three common suffixes: “ing,” “ion,” or “est”. Each word is then artificially split to two parts—the root word and the suffix. For example, the word “eating” is split into “eat” and “ing”, while “connection” is split into “connect” and “ion”. This ensures that processing the suffix token is necessary to reconstruct the identity of the word, and reduces possible effects of strong distributional artifacts of token co-occurrence.
Using logit lens, we identify FFN layers where the update to the residual stream can decoded as the original single-token word. We then ablate these layers by zeroing out their updates to the residual stream. As a control, we ablate the same proportion (5%) of random FFN layers.
Our results, shown in Fig. 12, reveal that ablating the identified FFN layers lead to a sharp drop in word retrieval rates, effectively disrupting the detokenization process. In contrast, ablating random layers has little to no effect on retrieval accuracy. This indicates that the FFN layers play a critical role in reconstructing word representations rather than simply enhancing contextualization.
Figure 12: Comparison of retrieval rates using logit lens for suffix-split words under three conditions: ablation of identified FFN layers (blue line), regular retrieval (orange line), and control with random FFN layer ablation (green line). Ablation of critical layers causes a sharp drop in retrieval accuracy, highlighting their importance in detokenization.
Appendix F Comparision Between Logit Lens and Cosine Similarity
To validate our use of the logit lens, we repeat the artificial split-word experiment using cosine similarity. We use the same setting of artifical splits based on suffixes detailed above in App. E.
Figure 13: Comparison of retrieval accuracy for split-word experiments using either logit lens (blue line) and cosine similarity (orange line) across model layers.
For the logit lens, we adapt the standard implementation to measure similarity between the hidden representation of the final token and the input embedding space (vocabulary space), rather than the output embedding space typically used to predict the next token. This adjustment allows us to directly evaluate the alignment between the hidden states and the embeddings of the original words in the vocabulary. In parallel, cosine similarity is used to compute the similarity between the same hidden states and the embeddings of the original words in the vocabulary space.
Our results, shown in Fig. 13, demonstrate that both methods produce nearly identical patterns. In both cases, retrieval accuracy peaks in the middle layers, where the hidden representation of the suffix token aligns most closely with the original word. Beyond these middle layers, retrieval accuracy declines, likely reflecting the model’s shift toward representing the prediction of the next token rather than maintaining the full representation of the current word.
These findings reinforce the validity of using the logit lens when adapted to the input embedding space, as a streamlined approach for analyzing model representations. This approach results in similar trends to cosine similarity while offering a more interpretable and direct framework for measuring alignment with the vocabulary space.
Appendix G Learning the Linear Maps $T_{\ell,E}$ and $T_{\ell,U}$
To expand the model’s vocabulary without modifying its core parameters, we construct linear transformations that map hidden states at different layers to the model’s embedding and unembedding spaces. By learning these transformations solely from the model’s existing vocabulary, we can infer new token representations without modifying any of the model’s core weights. This section details our method for learning these mappings.
G.1 Extracting Model Representations
Given a pretrained language model with input embedding matrix $E\in\mathbb{R}^{V\times d}$ and output unembedding (LM head) matrix $U\in\mathbb{R}^{V\times d}$, where $V$ is the vocabulary size and $d$ is the hidden dimension, we extract the representations used for learning the transformations as follows:
  • Embedding and Unembedding Matrices: We extract the model’s embedding matrix $E$ and LM head matrix $U$ before making any modifications to the vocabulary.
  • Hidden States Across Layers: For each token $t$ in the model’s original vocabulary, we pass $t$ as a single-token input and record its hidden state at every layer of the model. Let $h_{\ell}(t)$ denote the hidden state of token $t$ at layer $\ell$.
G.2 Learning the Linear Mappings
For each layer $\ell$, we aim to learn two linear transformations:
  • $T_{\ell,E}$ that maps hidden states to input embeddings.
  • $T_{\ell,U}$ that maps hidden states to unembedding representations.
We learn these mappings as orthogonal Procrustes problems (Schönemann, 1966), which seek to find the best orthogonal transformation aligning two sets of vectors. Specifically, for each layer $\ell$, we solve:
[…]
where $e_{t}$ and $u_{t}$ are the embedding and unembedding vectors of token $t$, respectively. The constraints enforce that each transformation preserves distances and does not distort the structure of the space. In our experiments, we use the Python implementation of Meng et al. (2022a).
G.3 Normalization with RMS Scaling
To preserve the relative magnitudes of embedding and unembedding entries, we normalize all representations using their root mean square (RMS) norm (Zhang & Sennrich, 2019). Specifically:
  1. 1.
    Preprocessing Training Representations: Before fitting the Procrustes transformations, we normalize all hidden states, embeddings, and unembedding vectors by dividing each vector $x$ by its RMS norm:
    $x_{\text{norm}}=\frac{x}{\|x\|_{\text{RMS}}},\quad\text{where }\|x\|_{\text{RMS}}=\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}.$ (3)
  2. 2.
    Applying the Learned Maps: When using the trained transformations on new detokenized representations $r$, we apply the following steps:
    1. (a)
      Normalize $r$ by its RMS norm: $r_{\text{norm}}=\frac{r}{\|r\|_{\text{RMS}}}$.
    2. (b)
      Apply the learned transformation: $\hat{e}=T_{\ell,E}r_{\text{norm}}$, $\hat{u}=T_{\ell,U}r_{\text{norm}}$.
    3. (c)
      Rescale by the mean RMS of the target space:
      $e=\hat{e}\cdot\mathbb{E}[\|e_{t}\|_{\text{RMS}}],\quad u=\hat{u}\cdot\mathbb{E}[\|u_{t}\|_{\text{RMS}}].$ (4)
This ensures that the new embeddings and unembeddings maintain a scale consistent with the original vocabulary.
Appendix H Efficiency Gains from Vocabulary Expansion
Beyond improving model performance on newly added words, our vocabulary expansion method directly reduces the number of tokens required to encode input text, and could lead to further potential efficiency gains in inference.
Dataset # Attempted # New Words Token Reduction
WikiText-103 14.1k 10.1k 10.5%
PubMed 9.5k 5.4k 13.5%
Wiki40B-Arabic 4.4k 0.7k 14.5%
Table 5: Reduction in average sequence length when encoding text with the expanded vocabulary for Llama2-7B. # Attempted is the number of multi-token words tested for expansion, while # New Words is those detected as detokenized using Patchscopes.
Table 5 summarizes the average reductions in sequence length achieved by encoding texts from each domain using the expanded vocabulary instead of the original vocabulary in the three datasets. Using our method, we attempt to expand Llama2-7B’s vocabulary with all multi-token words appearing at least $m$ times in the test set, where we use $m=1,5,50$ for WikiText-103, PubMed, and Arabic Wiki40B, respectively.
We find that the token reduction rates depend on the success rate of extracting detokenized representations using Patchscopes (which in turn depend on the textual domain and language). WikiText-103 achieves a 10.5% reduction in token count, with 72.9% of attempted words successfully converted into new vocabulary entries. PubMed shows a higher token savings rate (13.5%) despite a lower success rate (58.3%), as its vocabulary expansion targets more domain-specific multi-token terms. The largest efficiency gain is observed in Arabic Wiki40B, where encoding with the expanded vocabulary reduces token count by 14.5%, highlighting the method’s potential for languages with inherently longer token sequences. However, the success rate of identifying detokenized representations in Wiki40B is lower (16.0%), suggesting room for improvement in expanding vocabularies for morphologically rich languages, and suggesting many Arabic words are not stored in the model’s inner lexicon.
These results suggest that post-hoc vocabulary expansion can significantly reduce the computational cost of inference, particularly for non-English languages and domain-specific texts, without requiring any modifications to the model’s core parameters.